OpenStudy (anonymous):

Q: A lady buys goods worth rupees 200 from a shop.(shopkeeper selling th egoods with zero profit).The lady gives him 1000 rupees note.The shopkeeper gets the change from the next shop and keeps 200 for himself and returns rupees 800 to the lady????

OpenStudy (anonymous):

so suddenly life became a videogame?

OpenStudy (anonymous):

later the shopkeeper of the next shop comes with the 1000 ruppes note saying duplicate and takes his money back.How much

OpenStudy (anonymous):

how much LOSS did the shopkeeper face??? answer plzz

OpenStudy (anonymous):

how bro?

OpenStudy (anonymous):

sorry actually he lost 800(returned)+200(goods)=1000rupees

OpenStudy (anonymous):

2000 = total loss?

OpenStudy (anonymous):

no, he returned the 1000 rs which he took from other shopkeeper.so no gain or loss. he gave 800 rs to lady and goods worth 200 rs. so total loss=800+200=1000 rs.